alones (alones):

The graph of a line passes through the two points (-2, 1) and (2, 1). :/ I'M actually stuck here!!

pooja195 (pooja195):

https://www.desmos.com/calculator

alones (alones):

ummm wait what do i suppose to do there??

pooja195 (pooja195):

Type in your points it gives you the line

alones (alones):

well i don't need a line (thx by the way) i actually need to put it in a genral form ..the first time i got it x - y + 1 = 0but my sis says it's wrong :/

pooja195 (pooja195):

Im not fimiliar with that @TheSmartOne

TheSmartOne (thesmartone):

The "General Form" of the equation of a straight line is: Ax + By + C = 0

TheSmartOne (thesmartone):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AloneS The graph of a line passes through the two points (-2, 1) and (2, 1). :/ I'M actually stuck here!! \(\color{#0cbb34}{\text{End of Quote}}\) First find the slope do you know the slope formula?

alones (alones):

i think i don't know the slope :/

alones (alones):

this time i've got x + y - 1= 0 am i close to the answer ? :/

TheSmartOne (thesmartone):

\(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{x_1},\color{red}{y_1})~and~(\color{green}{x_2},\color{blue}{y_2})\)

alones (alones):

i'm kinda confused ...EH I got 0,1

alones (alones):

nah i mean i got (1,)0

TheSmartOne (thesmartone):

what did you get as the slope?

alones (alones):

i've got 0 and 1

alones (alones):

NVM I GOT IT RIGHT THX :) THX

TheSmartOne (thesmartone):

no problem :)